OpenStudy (anonymous):

A father is 3 times as old as his son. In 14 years' time, he will be twice as old as his son. How old is the father now?

OpenStudy (anonymous):

x=present age of son y= present age of father 3x=y ..........(1) y+14=2(x+14) y+14=2x+28 .......(2) substituting 1 in 2 3x+14=2x+28 x=14 y=3(14)=42

OpenStudy (anonymous):

Thanks! :)